OpenStudy (mackenzie2013):

Solve for n. 8 - n = -4 n =

OpenStudy (anonymous):

8-n=-4 -8 -8 n=-12

OpenStudy (solomonzelman):

add 8 to both sides divide both sides by -1

OpenStudy (solomonzelman):

@~Jane_Doe~ you are not correct.

OpenStudy (anonymous):

sorry should be n=12

OpenStudy (anonymous):

lol I saw that and was like what?!

OpenStudy (solomonzelman):

Sorry subtract 8 from both sides, not add

OpenStudy (anonymous):

lol.. sorry..

OpenStudy (solomonzelman):

everyone gets mistakes, happens.